In 2001, Tammy bought a home with her parents for $300,000, (assuming Tammy has 1/3 ownership). The home was upgraded for new solar panels costing $30,000. Two years later, her father died and just last year, her mother passed away. Saddened, Tammy finally moved out in January of this year. She plans to rent out the home for $3,000 per month, and had already spent $7,000 painting the house, inside and outside.

Tammy called and wanted to let you know her new situation. Her real estate agent advised her that the home was worth $570,000 when her father died, and $450,000 when her mother died. Given the fact that Tammy lives in a community property state, as her CPA, how would you advise her?
Answer #1

Community Property State is where the money earned by both the parties to the marriage and they invest it in a property with those earnings and such a property belongs to both Husband and Wife. Each spouse has half the right of the property. A spouse has the sole right to dispose of their separate property. A deceased spouse can distribute both their separate property and their share of the community property in a will.  

Under a community property state, when parents pass away, the Children does not get inheritance of the property. They get only the portion that they have ownership.

Total property value = $450000 (The latest computed value of the property has to be considered for sale)

Tammy's part = 1/3 * 450000 = $150000

Spendings : Solar panels = $30000

Painting = $7000

Sale value could be a minimum of = $113000

If Tammy decides to rent out the property she could only get 1/3 rd of the portion of the rent value

Rent value = 3000 * 12 = $36000 *1/3 = $12000 per year.

Hence it is advisable that Tammy sells her 1/3rd of portion of the property at the best selling price.